# CaloQVAE : Simulating high-energy particle-calorimeter interactions using hybrid quantum-classical generative models

### v1

@@ Abstract
This work is the first demonstration of the application of a quantum annealing device to the computationally expensive simulation of particle showers at the Large Hadron Collider. We have shown that it is possible to utilize the Quantum Processing Unit (D-Wave Chimera 2000Q) to generate RBM samples which can be used to generate particle showers. The resulting generated simulation is able to largely match the properties of the training dataset, and match the performance of other generative models. The raw QPU annealing time per sample is measured to be 20 $\mathrm{\mu s}$, which is 20 times faster than generating samples via GPU. While further work remains in evaluating on more realistic simulation and a wider range of physics topologies, the initial results are demonstrate CaloQVAE to be a promising application of quantum computing to open research questions in fundamental physics. The experimental results and code are available online [27].

### v3

@@ 5 Conclusion
This work is the first demonstration of the application of a quantum annealing device to the computationally expensive simulation of particle showers at the Large Hadron Collider. We have shown that it is possible to utilize the Quantum Processing Unit (D-Wave Chimera 2000Q) to generate RBM samples which can be used to generate particle showers. The resulting generated simulation is able to largely match the properties of the training dataset, and match the performance of other generative models. The raw QPU annealing time per sample is measured to be 20 $\mathrm{\mu s}$, which is 20 times faster than generating samples via GPU. While further work remains in evaluating on more realistic simulation and a wider range of physics topologies, the initial results are demonstrate CaloQVAE to be a promising application of quantum computing to open research questions in fundamental physics.
@@ Acknowledgements.
Data Availability Statement The analytical results and code are available online githubQVAE .